# PLLaVA : Parameter-free LLaVA Extension from Images to Videos for Video Dense Captioning

### v1

@@ Abstract
Vision-language pre-training has significantly elevated performance across a wide range of image-language applications. Yet, the pre-training process for video-related tasks demands exceptionally large computational and data resources, which hinders the progress of video-language models. This paper investigates a straightforward, highly efficient, and resource-light approach to adapting an existing image-language pre-trained model for dense video understanding. Our preliminary experiments reveal that directly fine-tuning pre-trained image-language models with multiple frames as inputs on video datasets leads to performance saturation or even a drop. Our further investigation reveals that it is largely attributed to the bias of learned high-norm visual features. Motivated by this finding, we propose a simple but effective pooling strategy to smooth the feature distribution along the temporal dimension and thus reduce the dominant impacts from the extreme features. The new model is termed Pooling LLaVA, or PLLaVA in short. PLLaVA achieves new state-of-the-art performance on modern benchmark datasets for both video question-answer and captioning tasks. Notably, on the recent popular Video ChatGPT benchmark, PLLaVA achieves a score of 3.48 out of 5 on average of five evaluated dimensions, exceeding the previous SOTA results from GPT4V (IG-VLM) by 9%. On the latest multi-choice benchmark MVBench, PLLaVA achieves 58.1% accuracy on average across 20 sub-tasks, 14.5% higher than GPT4V (IG-VLM). Code is available at https://github.com/magic-research/PLLaVA.
@@ 2 Related Works
Video Multi-modality Models process video input and generate responses according to user commands. Commonly, they incorporate a projection network [30, 22, 19], inter-modality attention [17, 18] or a modality perceiver [45, 32, 12] as learnable interfaces. These interfaces are instrumental in melding the spatial-temporal dynamics of videos with large language models’ (LLMs) processing capabilities, by transforming video content into a sequence of tokens that LLMs can adeptly analyze. BLIP [16] marked a significant milestone by integrating a frozen vision encoder with BLIP to enhance video processing efficiency, with only the newly added Q-Former learnable. Demonstrating remarkable zero-shot capabilities in Video Question Answering (VQA), it outperformed existing techniques of its time. Extending the innovations of its predecessors, Video-ChatGPT [30] introduced the trailblazing approach of video instruction tuning, along with creating a dataset of high-quality instructional data. This initiative set a new standard for assessing models through video-based text generation benchmarks. VideoChat [17] employed cross-attention mechanisms to skillfully condense video tokens, aligning user queries with the dialogue context to enhance the model’s interpretative capabilities. Building upon these advances, VideoChat2 [18] refined the approach with a multi-stage bootstrapping technique that honed in on modality alignment and instruction tuning, amassing a robust collection of high-quality video data for fine-tuning instruction-driven tasks. VILA [23] proposes more advanced training recipes. Further integrating modalities, Video-LLaVA [22] leveraged a pre-aligned encoder adaptable to both images and videos, facilitating shared projections and enabling synergistic training across image and video-related tasks. CAT [41] introduces both video and audio to futher enhance understanding.
@@ 3.1 Failure Cases Analysis for Applying Image MLLMs
Figure 4: When training with more samples, (a) and (b) show performance of n-frame method is stagnant while pooling can keep rising. Curves in (b) indicate the vulnerability of n-frame to prompt design while pooling is more robust.
@@ 3.3 PLLaVA
Figure 6: The framework of PLLaVA begins with processing a video from the user through ViT-L and MM projector, yielding visual features with shape $(T,w,h,d)$. These features undergo Adaptive Pooling, which effectively reduces both temporal and spatial dimensions. The pooled features are then flattened and concatenated with question embeddings, serving as input to the image Large Language Model to generate response to the user. The weights of the image LLMs are fused with LoRA weight learned under video samples.
@@
We formalize the pooling process for video features as follows, a model structure is shown in  Figure 6. After feeding video frames $\mathbf{X}\in\mathbb{R}^{T\times C\times W\times H}$ into the CLIP-ViT model and the multimodal projector, we obtain an encoded vision feature $X_{v}\in\mathbb{R}^{T\times w\times h\times d}$ for a video input. This feature is then passed through a parameter-free Adaptive Structure Pooling module and reduced to a smaller size. Given the desired feature dimension $T^{\prime}\times w^{\prime}\times h^{\prime}$, the process is formulated as:
@@
These features are then concatenated with text input embeddings and fed into the LLM to generate responses. We also include a LoRA [9] module to adapt the LLM to video-related generation tasks. In conclusion, the weights that will be updated are Multimodal Projector and LLM LoRA.
@@ 4.1 Experiment Setting
We leverage instructional video-to-text datasets to extend the capabilities of image MLLMs to handle video inputs. The training data are sourced from the VideoChat2 dataset [18], which embraces data for various video understanding tasks, including 27k for conversation combined from VideoChat [17] and Video-ChatGPT [30], 80k of classification from Kinetics [13] and SthSthV2 [6], 450k caption data from Webvid [2], YouCook2 [47], TextVR [37] and VideoChat, 117 reasoning data from NextQA [38] and CLEVRER [42] and 109K questioning answering data annotated from Webvid, TGIF [20] and Ego4D [7], totaling 783k instructional tuning data.
We evaluate our trained models with the following video-to-text benchmarks. First, the open-ended Video Question Answer (VideoQA) includes MSVD-QA [39], MSRVTT-QA [39], ActivityQA [43], and TGIF QA [20]. Responses in these question-answering benchmarks typically consist of single-word answers. GPT-3.5 is used to evaluate the accuracy (Accuracy, with answers true/false) and quality (Score, ranging from 0 to 5) of the models’ responses. Additionally, we adopt the Video-based Generative Performance benchmark (referred to as VCG Score), introduced by VideoChatGPT [30]. These benchmarks often involve longer answers, encompassing five aspects of video understanding: CI (Correctness of Information), DO (Detail Orientation), CU (Context Understanding), TU (Temporal Understanding), and CO (Consistency). The generation is also assessed using the GPT-3.5 model. Furthermore, we also use the multi-choice Question Answering benchmark, MVBench [18], comprising 20 tasks that demand nuanced temporal comprehension of videos. This benchmark does not necessitate evaluation from the GPT-3.5 model.
@@
PLLaVA is constructed upon the image MLLMs, LLaVA Next [26, 25] models 7B, 13B, and 34B. We utilize their pre-trained weights available in the Hugging Face library [^1] https://huggingface.co/docs/transformers/en/model_doc/llava_next and integrate an adaptive pooling module to reduce feature dimensions before passing the input visual features to the LLM generation component. For the pooling layer, we uniformly select 16 frames as input and set the target pooling shape to be $16\times 12\times 12\times d$, where $d$ corresponds to the input dimension of the LLMs. During training, we employ a batch size of 128 and a learning rate of 2e-5, with a cosine scheduler and a warmup ratio of 0.03. All the reported results are evaluated on models trained for 6250 steps. For evaluation, we adopt the GPT-3.5-turbo-0125 model across all the benchmarks.
@@ 4.5 Case Studies
(a) Badminton match.
(b) Street Saxophone.

### v2

@@ Abstract
Vision-language pre-training has significantly elevated performance across a wide range of image-language applications. Yet, the pre-training process for video-related tasks demands exceptionally large computational and data resources, which hinders the progress of video-language models. This paper investigates a straightforward, highly efficient, and resource-light approach to adapting an existing image-language pre-trained model for dense video understanding. Our preliminary experiments reveal that directly fine-tuning pre-trained image-language models with multiple frames as inputs on video datasets leads to performance saturation or even a drop. Our further investigation reveals that it is largely attributed to the bias of learned high-norm visual features. Motivated by this finding, we propose a simple but effective pooling strategy to smooth the feature distribution along the temporal dimension and thus reduce the dominant impacts from the extreme features. The new model is termed Pooling LLaVA, or PLLaVA in short. PLLaVA achieves new state-of-the-art performance on modern benchmark datasets for both video question-answer and captioning tasks. Notably, on the recent popular Video ChatGPT benchmark, PLLaVA achieves a score of 3.48 out of 5 on average of five evaluated dimensions, exceeding the previous SOTA results from GPT4V (IG-VLM) by 9%. On the latest multi-choice benchmark MVBench, PLLaVA achieves 58.1% accuracy on average across 20 sub-tasks, 14.5% higher than GPT4V (IG-VLM). Code is available at https://pllava.github.io.
@@ 2 Related Works
Video Multi-modality Models process video input and generate responses according to user commands. Commonly, they incorporate a projection network [33, 25, 21], inter-modality attention [19, 20] or a modality perceiver [50, 37, 14] as learnable interfaces. These interfaces are instrumental in melding the spatial-temporal dynamics of videos with large language models’ (LLMs) processing capabilities [40, 36, 5], by transforming video content into a sequence of tokens that LLMs can adeptly analyze. Parameter efficient learning schemes [11, 23, 7] are adapted to reduce the computational cost. Among them, BLIP [18] marked a significant milestone by integrating a frozen vision encoder with BLIP to enhance video processing efficiency, with only the newly added Q-Former learnable. Demonstrating remarkable zero-shot capabilities in Video Question Answering (VQA), it outperformed existing techniques of its time. Extending the innovations of its predecessors, Video-ChatGPT [33] introduced the trailblazing approach of video instruction tuning, along with creating a dataset of high-quality instructional data. This initiative set a new standard for assessing models through video-based text generation benchmarks. VideoChat [19] employed cross-attention mechanisms to skillfully condense video tokens, aligning user queries with the dialogue context to enhance the model’s interpretative capabilities. Building upon these advances, VideoChat2 [20] refined the approach with a multi-stage bootstrapping technique that honed in on modality alignment and instruction tuning, amassing a robust collection of high-quality video data for fine-tuning instruction-driven tasks. VILA [26] proposes more advanced training recipes. Further integrating modalities, Video-LLaVA [25] leveraged a pre-aligned encoder adaptable to both images and videos, facilitating shared projections and enabling synergistic training across image and video-related tasks. CAT [46] introduces both video and audio to futher enhance understanding.
@@ 3.1 Failure Cases Analysis for Applying Image MLLMs
Figure 4: Validation curve of baseline model and PLLaVA. When training with more samples, (a) shows the validation curve with in-distribution (IND) prompt and (b) show the validation curve with out-of-disribution (OOD) prompts. It is observed that n-frame method saturates fast and the performance even drops when training with longer time. is stagnant while pooling can keep rising. Differently, the performance of PLLaVA keeps increasing during the training and saturates at the end of training.
@@ 3.3 PLLaVA
Figure 6: The framework of PLLaVA begins with processing a video from the user through ViT-L and MM projector, yielding visual features with shape $(T,w,h,d)$. These features undergo average pooling, which effectively reduces both temporal and spatial dimensions. The pooled features are then flattened and concatenated with question embeddings, serving as input to the image Large Language Model to generate response to the user. The weights of the image LLMs are fused with LoRA weight learned under video samples.
@@
We formalize the pooling process for video features as follows, a model structure is shown in  Figure 6. After feeding video frames $\mathbf{X}\in\mathbb{R}^{T\times C\times W\times H}$ into the CLIP-ViT model and the multimodal projector, we obtain an encoded vision feature $X_{v}\in\mathbb{R}^{T\times w\times h\times d}$ for a video input. This feature is then passed through a parameter-free Adaptive Average Structure Pooling module and reduced to a smaller size. Given the desired feature dimension $T^{\prime}\times w^{\prime}\times h^{\prime}$, the process is formulated as:
@@
These features are then concatenated with text input embeddings and fed into the LLM to generate responses. We also include a LoRA [11] module to adapt the LLM to video-related generation tasks. In conclusion, the trainable weights include Multimodal Projector and LLM LoRA.
@@ 4.1 Experiment Setting
We leverage instructional video-to-text datasets to extend the capabilities of image MLLMs to handle video inputs. The training data are sourced from the dataset used in VideoChat2 [20], which embraces data for various video understanding tasks, including 27k conversation videos from VideoChat [19] and Video-ChatGPT [33], 80k data of classification tasks from Kinetics [15] and SthSthV2 [8], 450k captioned data from Webvid [2], YouCook2 [52], TextVR [42] and VideoChat, 117 reasoning data from NextQA [43] and CLEVRER [47] and 109K annotated questioning answering data samples from Webvid, TGIF [22] and Ego4D [9]. In total, we use 783k instructional tuning data.
We evaluate our trained models with the following video-to-text benchmarks. First, the open-ended Video Question Answer (VideoQA) includes MSVD-QA [44], MSRVTT-QA [44], ActivityQA [48], and TGIF QA [22]. Responses in these question-answering benchmarks typically consist of single-word answers. GPT-3.5 [34] is used to evaluate the accuracy (Accuracy, with answers true/false) and quality (Score, ranging from 0 to 5) of the models’ responses. Additionally, we adopt the Video-based Generative Performance benchmark (referred to as VCG Score), introduced by VideoChatGPT [33]. These benchmarks often involve longer answers, encompassing five aspects of video understanding: CI (Correctness of Information), DO (Detail Orientation), CU (Context Understanding), TU (Temporal Understanding), and CO (Consistency). The generation is also assessed using the GPT-3.5 model. Furthermore, we also use the multi-choice Question Answering benchmark, MVBench [20], comprising 20 tasks that demand nuanced temporal comprehension of videos. This benchmark does not necessitate evaluation from the GPT-3.5 model.
@@
PLLaVA is constructed upon the image MLLMs, LLaVA Next [29, 28] models 7B, 13B, and 34B. We utilize their pre-trained weights available in the Hugging Face library [^1] https://huggingface.co/docs/transformers/en/model_doc/llava_next and integrate an average pooling to reduce feature dimensions before passing the input visual features to the LLM generation component. For the pooling layer, we uniformly sample 16 frames as input and set the target pooling shape to be $16\times 12\times 12\times d$, where $d$ corresponds to the input dimension of the LLMs. During training, we employ a batch size of 128 and a learning rate of 2e-5, with a cosine scheduler and a warmup ratio of 0.03. All the reported results are evaluated on models trained for 6250 steps. For evaluation, we adopt the GPT-3.5-turbo-0125 model across all the benchmarks.
@@ 4.5 Case Studies
(a) Street Saxophone.
(b) Badminton match.